5. X-rays hit a lead block with an intensity of 200Wm-2. Lead has a linear attenuation coefficient of 27 cm-1. Calculate the intensity of the X-rays 15mm into the lead block.


1
Expert's answer
2021-03-26T11:20:06-0400

The intensity "I" of the beam at the depth "x = 15mm = 1.5cm" is given by the following equation (see https://www.spectroscopyonline.com/view/tutorial-attenuation-x-rays-matter):


"I = I_0e^{-\mu x}"

where "I_0 = 200W\/m^2" is the initial intensity, "\mu = 27cm^{-1}" is the linear attenuation coefficient. Thus, obtain:


"I = 200\cdot e^{-27\cdot 15} \approx 5.2\times 10^{16}W/m^2"

Answer. "5.2\times 10^{16}W/m^2".